OpenStudy (jackfrost33):

help please how would you solve for x if 3:12 x:16

OpenStudy (triciaal):



OpenStudy (calculusxy):

\[\large \color {purple}{\frac{ 3 }{ 12 } = \frac{ x }{ 16 }}\] This is a proportion. So you need to "cross multiply" to get \( \color {blue}{3(16) = 12(x)}\)

OpenStudy (calculusxy):

Simply and then solve for \(\color{blue}{x}\).

OpenStudy (triciaal):